Fermi National Laboratory
 

Accelerator Update for October 5, 2007 - October 12, 2007

 

Fermilab Shutdown Report
10/5/07 – 10/12/07

Operations

Current status

  • Linac – operational, NTF treatment
  • Booster – commissioning
  • MI – Controlled Access
  • MI-8 – no access
  • Pbar Rings – controlled access
  • Transport – controlled access
  • A-E – Controlled access
  • F-sector – controlled access
  • Transfer Hall – controlled access

Run Coordinator

  • Tevatron Today
    • Continued work at A1 (Cryo)
    • E1 work continues with a leak repair at E16. Current estimates would have the house turned over to Cryo Dept. by late today
    • Find GF in A-E/F-sector
    • Verify/Clean Dust in F-Sector (P3 Line Magnets). There are other areas in TH/A-E that need minor attention.
    • EE Support TeV Elec. Walk Through
    • Later A1, F-Sector if dusting still in progress
    • E1 Walk through Saturday evening shift?
    • Secure CDF and D0 this afternoon.
    • Remaining Safety test?
  • Main Injector Today
    • Need access to 8GeV line (1-1.5hrs) Check Air leak for MI8 air compressor
    • This will take Booster Beam away
    • Verify LCW valves in MI/8GeV
    • Check New Collimator Motion controls in MI
    • No significant Mechanical items left
    • Find clear GF in MI10/MI60 area
    • Possibly Energize Kautz Road Substation Late day shift early evening shift.
      • If so, possibly ramp MI on Evening shift with Experts present
  • On Friday Eve We will-
    • Switch on Kautz Road Substation if ComEd is ready. If this can not be done by 1800hrs, will do it Saturday as planned at 0700-0730
    • Affected parties have been given a heads up on this possibility. MiniBooNE, NuMI, RF, Cryo, COntrols, Mechanical Support
    • Recover controls/vac/LCW etc.
    • Ops will verify ramps in Cards after restores are completed
    • Switch on MOS 86, 87, 89
    • Establish ramp with MI expert present in MCR
  • On Saturday We will-
    • Switch on Kautz Road Substation (if not done Fri eve)
    • Recover controls/vac/LCW etc. (if not done Fri eve)
    • Get MI RF on in the morning (4hrs)
    • Establish ramp if not done Friday eve
    • EE Support checkout of various MI/Beamline Supplies. Some are critical devices.
    • Complete various CDC Safety Tests as we have feeder 89 energized to turn on CDC PS
    • Establish MI 8 GeV beam.
    • Possibly run beam to MiniBooNE
  • Beam Permits for Operation
    • Today we will need if ready or possible
      • MI Beam Permit Sign-off
      • Pbar Beam Permit Sign-off
      • MiniBooNE Beam Permit Sign-off
      • Recycler Beam Permit Sign-off?
    • Monday
      • Tevatron Beam Permit Sign-off
  • Week 10 is Almost Done
    • Groups should Update call-in list today
    • We are in Commissioning mode
    • Ops are very busy with start-up
      • Please remember to coordinate all requests through Bob or Dan for operator assistance. If every group asks for a little thing it all ads up quick!
  • Thanks to Everyone for their hard work this shutdown
    • Many of the people that have worked so hard this shutdown will be called upon again during commissioning, Thanks in advance
  • We should continue to work carefully and take the time need to work safely during Start-up

LINAC

  • Current Linac Status
    • Linac Output is ~ 36.5 mA
    • Linac Loss Sum ~ 10 (Average)
  • Work in Progress
    • Linac Tune Up
      • L.E. Linac Beam Feedback Tune Up
      • L.E. Linac Modulator Tune up on 2 & 4
      • H.E. Linac Klystron Tune up
    • Terminate cables for LLRF Upgrade
    • Retune LRF4 & LRF5 Phase Lock Signals
    • Calibration & Treatment
      • M, Th, F (7:00 a.m. to 12:00 p.m.)
    • Dosimetry
      • Not scheduled next week
  • Weekly Status
    • 96% Availability
    • 6 Hours of down time
  • Major Downtime
    • LRF2 (6 hours) Keyer Driver Replacement
      • Also replaced 6544 tubes
    • Klystron Debuncher problems
    • LRF3 Station Trips on RF3CEN bit
  • Post-Shutdown Work Request
    • (4+ Hrs) Continue to tune LRF2 and LRF4 Modulator Regulator Loop
      • May be causing the energy variation in the Linac
      • Will take significant investment in downtime to try to improve beam loading
    • (2 Hrs) Continue to tune LRF5 after changing 9” transmission line coupling.
      • This will only minimally interfere with normal operations
      • Does not require dedicated study time
    • (2 Hrs) Switch over from Pulse Repeater System Start Trigger over to individual 15 Hz trigger from each system node
      • Compete system turn down and start up
      • Needs at least one day notice to set up timing channels
    • (40 Hrs over 5 days) Replace Turbine Flow meters on all klystron water skids
      • Can be done in parts and needs mechanical support
    • (4 Hrs) Test new LLRF: Proof of principle experiment on station 5
      • This will not interfere with NTF operation
      • Not currently prepared for this study

Booster

Booster Beam Startup

  • Beam Accelerated to 8 GeV Friday – 5th
  • Beam Accelerated to Booster Dump Monday - 8th (MI WATER BACK!!)
  • Booster RF 7 removed from tunnel (repaired put back on Wednesday)
  • Booster Bump Ratios worked on Monday – Thursday
  • Booster Orbit Work Thursday – Friday
  • Booster Efficiency 85% @ 3 Turns

Work yet to be Done

  • BLM at extraction (BLM026) not working – requires tunnel inspection.
  • Higher intensity – tuning of higher order ramps
  • Higher intensity tuning of ramps (corrector ramps are still disabled)
  • Extraction timing and tuning

Antiproton Source

Mechanical

  • A10 Obstruction Search & Shuffle
  • Install new skew quad
  • Water leak on IB6 manifold
  • IQ11 water hose connection
  • Deb DCCT fix

Target Station

  • 10mm-3 into transformer
    • Testing
  • Installation of samples for T972

Electrical (Pbar & EE Support)

  • Regulation Electronics
    • Done except A:LQ

Other

  • Interlocks/Safety System
    • Vault key tree done; testing being done
  • Repair cathodic protection
    • Schedule for 10/12
  • Holes under buildings
    • Water dripping from tunnel ceiling
      • Contractor sealed ceiling
    • Contractor filled holes
  • DCCT calibration
  • SEM maintenance
  • Clogged sump line – Transport

Draft Scheduled agenda

  • Today and Saturday run PS
  • Sunday afternoon/evening: establish reverse protons to Accumulator
  • Reverse protons to Accumulator & Debuncher
  • Place holders and dependent upon needs of the rest of the complex
    • Thursday: start stacking
    • Friday: transfers to Recycler

To Only Reestablish Operations

  • Reverse Protons
    • {1 shift} re-establish MI-ACC ($2D/30sec)
    • {3+ shifts} Accumulator (1-shot timeline)
      • BPM checkout; Orbit Checkout; Aperture (A10, running wave); Tunes across aperture; Calibrate emittance monitors
    • {1+ shifts} Debuncher (1-shot timeline)
      • Establish beam; Orbit Checkout; Aperture (EKIK, running wave)
  • Stacking
    • {1+ shifts} 120GeV operations ($29/4.4sec)
      • Protons to Target; Secondaries to Deb (include 120GeV orbit for lens height); Kickers & RF systems timing; antiprotons to core
    • {2 shifts} Check phasing of cooling systems
    • 20mA {<0.5 shift} phase dampers
    • 40mA to 60mA transfer to Recycler

Further Work

  • Clean Accumulator Studies [Reverse Protons]
    • {1 shift} D/A optics ($2D/min)
      • Deb Rev Prot TBT; D/A BPMs?; SEMs
    • {1+ shift} D/A Transfer Efficiency ($2D/min)
      • Set Acc transfer scrapers to define different beam sizes
        • Record DCCTs, SEMs, Loss monitors
    • {1+ shift} A10 & running wave (1 shot timeline)
      • Make component limiting aperture and perform running wave
        • Do different limiting apertures cause running wave response?
    • {1 shift} Accumulator optics – all orbits (1 shot timeline)
    • {1 shift} Commission SQ107 (1 shot timeline)
      • Includes coupling correction and measuring vertical dispersion
  • Accumulator hold onto core [Reverse Protons]
    • {1 shifts} Deb Optics (1 shot timeline)
      • Measure; Change; Measure
  • Stacking
    • {0.5 shift} Commission Debuncher switch-in filters
    • {1+ shift} D/A Transfer Efficiency (long cycle time; consistent beam)
      • Vary Debuncher cooling time
  • Record IPM, Loss Monitors, production rate

Main Injector

Beam Startup Activities

  • Power Supply Checkout
    • A1, P1, P2 and abort lines need EE support to turn on. (6 hrs.)
    • This cannot occur at the same time as Tevatron turn on
  • Critical device safety testing
  • Measurements requiring dedicated time
    • Establish 8 GeV beam. 4 hrs
    • MI8 line lattice measurements and matching 2 x 4 hrs
    • Establish 120 GeV beam. Smooth orbits4 hrs
    • Aperture scans at MI30 2 hrs.
    • Quad moves? 8 hrs
    • Coalescing tune up 4 hrs
    • Slip Stacking tune up. 2 hrs
    • Tune up extraction positions and abort line 4 hrs per line (6)
  • Friday
    • Establish ramp if possible (slow warm up)
  • Saturday
    • Turn on RF
    • Critical device tests
    • Establish 8 GeV beam, tune abort
    • Aperture scans at MI10, MI30, MI40, MI52
    • Accelerate to 120 / 150 GeV
  • Sunday
    • MI8 line lattice measurements (morning)
    • Extract 8 GeV beam to Accumulator
    • Measure orbits at high field
  • MI Status

Recycler

  • Finished Before this Week
    • New DCCT installed, baking finished
    • V303 and H308 moved for MI collimator work (done)
    • STC tank 7 leak check finished, nothing new
    • Stochastic cooling
      • New attenuators (done)
      • 1-2 GHz conversion (done) to H Betatron cooling
    • Bulk supply upgrade (done)
    • The 800 line leak repair (done)
    • Diagnostics maintenance
      • New FW paddles (done)
      • Schottky (1.75 GHz) system gain check (done)
      • Schottky (21 MHz) L.O. check (done)
    • Phase Trombone installation (done)
    • QCL cable positioning from 304 to 309 (done)
    • Magnet moves at 305 and 307 (done)
    • TOR330 repair (done)
    • Phase Trombone wiring finished
  • Since Last Week
    • BLM for scrapers replacement (done)
    • BPM calibration check and repair (done)
    • Elevation survey results
    • Field check for new quadrupoles and ring wide corrector check - done
    • Kicker timing verification
    • RF power and correction system check
    • Damper status report (next week)
    • Safety sign-off (finished for Pelletron study; one more for running Recycler beam)
    • MW803 stuck; reported to Instrumentation
    • End-of-shutdown check (done)
    • Recycler Start-up plan ready (next report)
    • Tunnel walk (done)

Tevatron

  • Vacuum Issues
    • E-18 conning tower had intermittent short Open up and fixed
    • A-1 certified and ready for Hi pot
    • All other houses read to go
  • Items Accomplished
    • Measured rolls all dipoles and quads
    • AMG measured and adjusted over 100 magnets around the Tevatron
    • D-16 quad un rolled
    • All ceramics cleaned
    • 2 spools and 2 dipoles replaced
    • 1 dipole moved into aisle and repaired
    • 6 houses warmed up and certified (only planned on doing 4 houses)
    • Conning towers at E1 repaired
    • C0 vacuum system rebuilt and baked out
    • HLS systems around the ring serviced
    • HLS systems on low beta quads rebuilt
    • Replaced F 17 kicker
    • Low beta quads at B0 and D0 realigned
  • CDF Low Beta Shielding

NuMI

  • MI65 emergency power transfer switch repair
    • New controller installed and tested
  • Target, Horn 1 & Horn 2 maintenance and upgrades
    • Changed water filters for Horn 2 RAW
    • Re-adjusted all the flows for Horn 1 RAW
    • Removed all four old module bushing blocks from Horn 2
    • Installed four new floating module bushing blocks on Horn 2
    • Hipot Horn 2 in work cell & in target chase
    • Installed floating water lines for Horn 2
  • Horn 2 suction line repair
    • Completed Horn 2 suction line repair on Sunday
    • Repair took approximately 2.5 hours
    • Total dose 343 mR
    • Presently checking for leaks
  • Decay pipe window and helium line installation
    • Continue to hold weekly coordination meetings
    • ODH analysis in process (ODH 0 likely)
    • Final drafts of Shielding Assessment Addendums complete
    • Draft of “Analysis of the NuMI DP Window for Operation with Helium Pressure” engineering note completed
    • Piping System & Low Pressure Vessel Engineering notes, in process
    • SAD update in process
    • Operational readiness review process in place
  • Target pile dehumidification installation project (85% Complete)
    • Continue to install 26” & 16” round duct on dehumidifiers
    • Completed installation of 1” SS condensate piping
    • Continue to install 4”chilled water piping in dehumidifier area
    • Installed evaporator vent pipe in SW corner of service building
    • Continue electrical installations in Target Hall and Service Building
    • Continue installation of chiller and pumps in Mezzanine
  • Seal MINOS Hall roof for Minerva experiment (50% Complete)

MiniBooNE

  • Need signatures:
    • Systems: Operations
    • Final Approvals: (all)

MuCool Beamline Installation

  • Mechanical work
    • Beam stop #2 assembled, tested, delivered and installed
    • All stands installed and secured
  • Alignment
    • Work continue with the referencing of MW system
  • Electrical work
    • Inst. group connected all BPM, BLM and toroid cables
  • Vacuum work
    • All vacuum connections are complete
    • Initial rough vacuum by 10/11. Milestone accomplished
      • No leaks were found at this time
      • Leak check is expected to take place today
        • In the near future, we’ll have to break vacuum in order to install the MW apparatus inside the vacuum cans
  • Other activities
    • MTA safety system testing complete
    • Access into the enclosure must be done under control access
  • The MTA beam line has been installed!
    • Many thanks to AD, PPD and TD for their support and for making their exceptional experts available to work on this project. Their enthusiasm and perspective was exactly what this project needed in order to succeed

Collider Detector at Fermilab

  • Plugs bolted down
    • Solenoid startup started
  • Flammable gas flow to all detectors
    • Wire chamber HV raised (most of it)
    • Wire chamber checkout continues
  • Much (and most) other work completed
  • 24/7 process system technician coverage
  • Owl shift on standby pending beam - note a configuration change
    • CDF TeV abort is now the logical OR of old BLM and new diamond systems
    • Still one signal coming to you
  • Subfloor secured yesterday afternoon, hall this afternoon
  • Detector elevation survey shows detector rising from last survey to this
  • We go all out for fire protection week; a fire drill wasn’t good enough for CDF…

D-Zero

Detector closed

Alignment

All Jobs Completed

  • Alignment and Metrology Home page

Mechanical

  • Need access to the 8 GeV line
  • The Switchyard 120 short was found
  • There is still MINOS helium work to do

Cryogenic System

Current Operations

  • Tevatron cooling down to liquid helium temperature except A1 and E1 at room temperature
    • A1 and E1 reconnection for subsequent cooldown in progress
    • Contamination plugging of FRIG F3, F4, and A3 heat exchangers being investigated
    • CHL Helium Coldbox-1 operating fine on compressors B and D.
    • 2 nd delivery of bulk liquid helium being delayed until contamination issue at FRIG buildings is resolved
  • CHL Nitrogen Reliquefier (NRL) plant online

Controls

People will have to restore systems after the power outage on Saturday

FESS

MSS & KRSS

  • MSS energized on 10/3/07.
  • KRSS maintenance WIP thru today. To be energized on sat 10/13/07. Work included transformers, buss ducts, air switches/fuses. No major findings. Yard was relamped

CUB

  • DDC work – Pbar 95 LCW flow control valves/actuators work-in-progress (WIP)
  • Pbar S/T heat exchanger [HE-8] is still under evaluation regarding end of life from Eddy current testing to be finalized next week which shows up to 30% of its tubes are over/at/near the 50% thickness factor [15 leaking tubes were plugged]. This is the back-up exchanger for the P/F HE-9. We will be able to continue to use but plans should be made for its replacement
  • The 55LCW pump replacements are complete [LP4 & 5]. LP5 is on line.
  • DI regenerations have been in progress causing tight scheduling for approved releases to Batavia sanitary
  • A new larger ICW strainer for make up is being installed within the next week

Casey’s Pond

  • ICW 16 inch FLOW METER/PIT WIP. End of month completion date, which should not impact on the experiments. The new enclosure also includes the modified Brominating system

Craft SD Status

  • Electrical craft jobs have been prioritized to finish critical work
  • Mechanical & HVAC work is completed
  • Any remaining work will be entered on the Electronic work list
  • FINDINGS – our PdM program for vibration analysis & electrical IR scanning on hot connections/devices was utilized along with our PM/ inspections, which turned up several corrective work orders (WOs)
  • We completed over 2,000 WOs & over 5,000 hrs of work

Other Jobs

  • TENTATIVE SITE ICW FLUSH was started Mon 10/01 & completed on Fri 10/05. There was minimal Zs throughout most of the sight except for the North Meson areas such as buildings MP9 to A – D. Switch Yard was ok
  • ZEBRA MUSSEL TREATMENT SCHEDULED from today starting at 7:00 am thru ~7:00 am Sat. 10/13. These treatments will be routinely scheduled every year to control growth
  • WILSON HALL FIRE PUMP RUN [ICW] test to be scheduled by end of October to satisfy DOE requirements
  • WILSON HALL power shutdown TBS in November
  • FESS OPERATIONS is open to any post SD discussions regarding the work we did & any improvements with scheduling & communications

Safety

All required interlock testing has been completed.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 10/12/2007    email Fermilab